OpenStudy (babyslapmafro):

Please help me with finding the derivative of the following problem... y=ln(tanx)

OpenStudy (babyslapmafro):

\[\frac{ dy }{ dx }=\frac{ 1 }{ tanx}\times \sec x\]

OpenStudy (babyslapmafro):

oops shouldbe sec squared x

OpenStudy (anonymous):

sec(x)^2

OpenStudy (babyslapmafro):

\[=\frac{ \sec ^{2}x }{ tanx }\]

OpenStudy (babyslapmafro):

answer in the book is 2csc2x

OpenStudy (anonymous):

sec squared x is the same as 1 over cos squared x and 1 over tan x is cos x over sin x so you end up with 1 over cos x sin x which is the same as a half sin 2x and 1 over a half sin 2x is 2 csc 2x

OpenStudy (anonymous):

sorry in the middle there i just wanted to say cos x sin x is a half sin 2x